How do you solve:
What is the slope of the line with the equation 2x + 3y + 6= 0?
Thank you

Answer by dlam5(15)

2x +3y +6 =0
Turn into slope-int form.
(y=mx + b; recall m=slope)
3y = -2x -6
/3 /3 /3
y= -2/3x - 2
m = -2/3(slope)